You’re trying to save to buy a new $240,000 Ferrari. You have $41,000 today that can be invested at your bank. The bank pays 4.8 percent annual interest on its accounts.
  
How long will it be before you have enough to buy the car? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

240,000=41000*(1.048)^n

(240,000/41000)=(1.048)^n

Taking log on both sides;

log (240,000/41000)=n*Iog 1.048

n=log (240,000/41000)/log 1.048

=37.69 years(Approx).